OpenStudy (anonymous):

Suppose that f(2) = 5, f '(2) = 5, g(2) = -6, and g'(2) = 4. Find the following values: (fg)'(2); (f/g)'(2); (g/f)'(2)

OpenStudy (anonymous):

write down the formulas, for example for the product rule, and then plug in the number you are given. for example \[(fg)'=f'g+g'f\]

OpenStudy (anonymous):

number would be \[5\times (-6)+4\times 5\]

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

And do the same for the other problems by using rules?

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

Perfect, thanks =]

OpenStudy (anonymous):

just write down the rules \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] and then plug in the values given for each

OpenStudy (anonymous):

yw